Aggregator
COMMENT ON C100-03-R Cryomodule Installation in 2L25
"North Linac: Beam Permit"
Autolog restore
Autolog save
Replaced camera for AA SLM
COMMENT ON 2L25 Cryomodule removal from linac
COMMENT ON Install 2L24, 2L26 utubes
Day 1 activities were accomplished nominally on 2/4. Day 2 activities were accomplished nominally on 2/5.
When the L-tubes were removed on 2/5 in preparation for the swap to the Primary Return u-tubes, the modules pressurized excessively, requiring back-out and putting the modules on a path to 300K warmup (and a repeat of the activities described in this ATLis).
https://logbooks.jlab.org/entry/4468629
https://logbooks.jlab.org/entry/4468629
iochdcol restored
COMMENT ON Replace Multiple Faulty 20A SCE
COMMENT ON Vacuum Activities CEBAF 2/18/2025 through 2/18/26
Nitrogen O2 read back
I&C was investigating the alarmed read back for the Tank O2 Level% to the nitrogen generator. The following measurements were taken after isolating the signal source from the generator. I&C injected signal into the system at 4ma establishing a 2V input to the ADC. Reflecting in a read back of 0 in Epics and clearing of the alarming status. I&C proceeded to inject 8, 12, 17, 20 ma in the the system. All values scaling and reflecting accurately in Epics. Verify all I&C hardware and software configurations.
COMMENT ON BSY ion pump swap
OPS-PR associated with the work for reference and later comment / closure: https://logbooks.jlab.org/entry/4353300
COMMENT ON BSY ion pump swap
IOC MAC address change
Autolog restore
Autolog save
Autolog restore
Autolog save
COMMENT ON MST magnet installation and related magnet work at 1L09,2L03, 2L25, 1L26
Mag Shop:
We just completed magnet work at 1L26/27, DC power can re-hookup back up when ready.
Hope to go down first thing in am and work @1L09, on new magnet (MST) install let us know soon as ready with call or email.
DC power notified as well as all other teams.










